OpenStudy (anonymous):

find the difference 3-(-9)

OpenStudy (mathstudent55):

-(-a) = a

OpenStudy (anonymous):

can someone help me

OpenStudy (mathstudent55):

-(-a) = a -(-9) = 9

OpenStudy (mathstudent55):

3 - (-9) = 3 + 9

OpenStudy (anonymous):

o thx u wanna help me with more

OpenStudy (mathstudent55):

Sure. Start a new post, I'll help you there.

OpenStudy (mathstudent55):

To subtract numbers, change the subtraction into an addition, then follow the rules of addition. To change a subtraction into an addition, change minus into plus and change the sign of the second number. Example 1: -3 - (-5) = -3 + 5 = 2 Example 2: 23 - 28 = 23 + (-28) = -5

OpenStudy (anonymous):

i did